# Unified Framework of Forced Magnetic Reconnection and Alfvén Resonance

### v1

@@ VII Summary and conclusions
In this paper we have derived and analyzed the general solution to the resistive, linearized MHD equations (Eqs. 4-5) describing the evolution in time of a plasma with a sheared magnetic field (Eq. 1) subjected to an oscillatory boundary perturbation turned on at time $t=0$ (Eqs. 8-9). Through asymptotic matching, we determined the spatial dependence (Eq. (28)) and the amplitude (Eq. (30)) of the general solution for the perturbed stream function in the inner layer. Such a solution shows explicitly that the key parameter to describe different regimes in the evolution of the system, as well as the transition from forced reconnection to Alfvén resonance, is $Q=\hat{g}S^{1/3}\hat{k}^{-2/3}$. We have then discussed two asymptotic limits of Eq. (28), $Q\gg 1$ and $Q\ll 1$, and we demonstrated that our theory describes Alfvén resonance when the frequency of the forcing is $\hat{\omega}_{0}\gg S^{-1/3}$, and forced reconnection for small to zero frequency, $\hat{\omega}_{0}\ll S^{-1/3}$. Spatio-temporal solutions have been derived for a standing wave boundary condition, given in Eq. (33), which we compared with linear simulations.
@@
In the ideal limit, defined by $Q\gg 1$ and $z/Q^{3/2}\ll 1$ (up to first order), our general solution recovers that of forced reconnection for $\hat{\omega}_{0}=0$; for $\hat{\omega}_{0}\neq 0$, our solution describes Alfvén resonance. In particular, the known logarithmic singularity at the resonance point where $\omega_{0}/k=v_{a}(x)$ is due to the residues at the poles of the boundary driver, and contribute to the time-asymptotic response of the system, as can be seen from the Bromwich contour in Figure 1.
@@
The non-ideal time-asymptotic response of the system corresponds to $Q\ll 1$. We have provided the corresponding asymptotic expansion of the inner layer stream function for $Q\ll 1$, by recovering the constant-psi solution that describes forced reconnection for times $\hat{t}\gg S^{1/3}$. The corresponding Bromwich contour is shown in Figure 2.
@@
For large frequencies, the time-asymptotic response is dominated by the residues at the forcing frequency (see Fig. 2), and if the condition $\hat{\omega}_{0}S^{1/3}k^{-2/3}\gg 1$ is satisfied, then a different expansion of the stream function for their evaluation is required. In this case, the solution is given by the asymptotic expansion of the inner layer solution for $Q\gg 1$ with terms in $z/Q^{3/2}\ll 1$ retained to higher order. This regime represents Alfvén resonance in the time-asymptotic limit.
@@
A comparison of the theoretical predictions and numerical simulations is given in Figure 3 and 4, testing our model’s ability to capture both forced reconnection and Alfvén resonance. We show that for large frequencies the resonant poles of the boundary driver were sufficiently far from zero (accordingly, the resonant points are far from the neutral point) and a pair of anti-symmetric resonances, of width $\delta\sim S^{-1/3}$, are clearly defined in that case. As these poles are brought towards $\hat{\omega}_{0}=0$ (the resonant points approach the neutral point $\hat{x}=0$) in the small frequency limit, the widths of the resonant layers overlap across $\hat{x}=0$ by coupling to the reconnection mode. The rate of reconnection and its signatures are strongest when the poles overlap entirely, that is for $\hat{\omega}_{0}=0$.
@@
Finally, we demonstrated that our solution recovers known scalings of the resonant layer $\delta$ as a function of the Lundquist number $S$ under specific limits. Additionally, we demonstrated new scalings of the reconnection rate and reconnected flux as functions of the driving frequency of the boundary condition. In Figure 5 we show that, in the constant-psi regime, the maximum reconnection rate and reconnected flux are largest for frequencies less than $\hat{\omega}_{0}\lesssim S^{-3/5}$ and is reduced at larger frequencies until reconnection is completely suppressed for $\hat{\omega}_{0}\gg S^{-1/3}$.
@@
We conclude that forced reconnection is indeed a particular case of Alfvén resonant absorption and there is a transition from Alfvén resonance to forced reconnection when the forcing frequency $\hat{\omega}_{0}\ll S^{-1/3}$. However, the reconnection rate and reconnected flux are strongly reduced for $\hat{\omega}_{0}\gg S^{-3/5}$. When the forcing frequency is $\hat{\omega}_{0}\gg S^{-1/3}$, Alfvén resonance decouples from the reconnecting mode and the formation of resonant layers away from the neutral line effectively shields reconnection.
@@ Appendix A Numerical Simulations
$\frac{\partial\psi}{\partial\hat{t}}=\hat{x}\frac{d\phi}{d\hat{y}}+\frac{1}{S}\hat{\nabla}^{2}\psi$ (55)
$\frac{\partial\Omega}{\partial\hat{t}}=-\hat{x}\frac{d}{d\hat{y}}\hat{\nabla}^{2}\psi$ (56)
where $\Omega=-\nabla^{2}\phi$ is the vorticity. We impose a boundary condition at $\hat{x}=\pm 1$ with small amplitude given by:
$\psi(\hat{x}=\pm 1,\hat{y},\hat{t})=\Xi_{0}\cos(\hat{k}_{0}\hat{y})\cos(\hat{\omega}_{0}\hat{t})\left[1-\left(1+\frac{\hat{t}}{\tau}\right)e^{-\hat{t}/\tau}\right],$ (57)
$\phi(\hat{x}=\pm 1,\hat{y},\hat{t})=\pm\Xi_{0}\frac{\sin(\hat{k}_{0}\hat{y})}{\hat{k}_{0}}\left\{\left(\frac{\hat{t}}{\tau}\right)^{2}e^{-\hat{t}/\tau}\cos(\hat{\omega}_{0}\hat{t})-\hat{\omega}_{0}\sin(\hat{\omega}_{0}\hat{t})\left[1-\left(1+\frac{\hat{t}}{\tau}\right)e^{-\hat{t}/\tau}\right]\right\},$ (58)
with $\tau$ being a small growth parameter, following [3]. The code is periodic in the $y$ direction, where the Fast Fourier Transform is used to calculate spatial derivatives. The $x$ direction is non-periodic and a finite difference scheme of the sixth order is used to calculate spatial derivatives [27]. We use an explicit Runge-Kutta of the third order to advance in time. We impose the 2/3 rule for dealiasing with respect to the periodic variable, and a pseudo-spectral filter in $x$ [27]. Simulations are performed by using a grid $L_{x}\times L_{y}=[-1,1]\times[0,2\pi]$ with mesh points $N_{x}\times N_{y}=512\times 32$ for $S=10^{4}$, $N_{x}\times N_{y}=1024\times 32$ for $S=10^{5}$, $N_{x}\times N_{y}=2048\times 16$ for $S=10^{6}$, and $N_{x}\times N_{y}=4096\times 16$ for $S=10^{7}$, with $N_{x}$ chosen to resolve their respective boundary layers. In all simulations we have fixed $\tau=0.001$, $\Xi_{0}=0.001$, and wavenumber $\hat{k}_{0}=1$ while varying $S$ and the frequency $\hat{\omega}_{0}$.

### v2

@@ VII Summary and conclusions
In this paper we have derived and analyzed the general solution to the resistive, linearized MHD equations (Eqs. 4-5) describing the evolution in time of a plasma with a sheared magnetic field (Eq. 1) subjected to an oscillatory boundary perturbation turned on at time $t=0$ (Eqs. 8-9). Through asymptotic matching, we determined the spatial dependence (Eq. (28)) and the amplitude (Eq. (30)) of the general solution for the perturbed stream function in the inner layer. Such a solution shows explicitly that the key parameter to describe different regimes in the evolution of the system, as well as the transition from forced reconnection to Alfvén resonance, is $Q=gS^{1/3}k^{-2/3}$. We have then discussed two asymptotic limits of Eq. (28), $Q\gg 1$ and $Q\ll 1$, and we demonstrated that our theory describes Alfvén resonance when the frequency of the forcing is $\omega_{0}\gg S^{-1/3}$, and forced reconnection for small to zero frequency, $\omega_{0}\ll S^{-1/3}$. Spatio-temporal solutions have been derived for a standing wave boundary condition, given in Eq. (33), which we compared with linear simulations.
@@
In the ideal limit, defined by $Q\gg 1$ and $z/Q^{3/2}\ll 1$ (up to first order), our general solution recovers that of forced reconnection for $\omega_{0}=0$; for $\omega_{0}\neq 0$, our solution describes Alfvén resonance. In particular, the known logarithmic singularity at the resonance point where $\omega_{0}/k=v_{a}(x)$ is due to the residues at the poles of the boundary driver, and contribute to the time-asymptotic response of the system, as can be seen from the Bromwich contour in Figure 1.
@@
The non-ideal time-asymptotic response of the system corresponds to $Q\ll 1$. We have provided the corresponding asymptotic expansion of the inner layer stream function for $Q\ll 1$, by recovering the constant-psi solution that describes forced reconnection for times $t\gg S^{1/3}$. The corresponding Bromwich contour is shown in Figure 2.
@@
For large frequencies, the time-asymptotic response is dominated by the residues at the forcing frequency (see Fig. 2), and if the condition $\omega_{0}S^{1/3}k^{-2/3}\gg 1$ is satisfied, then a different expansion of the stream function for their evaluation is required. In this case, the solution is given by the asymptotic expansion of the inner layer solution for $Q\gg 1$ with terms in $z/Q^{3/2}\ll 1$ retained to higher order. This regime represents Alfvén resonance in the time-asymptotic limit.
@@
A comparison of the theoretical predictions and numerical simulations is given in Figure 3 and 4, testing our model’s ability to capture both forced reconnection and Alfvén resonance. We show that for large frequencies the resonant poles of the boundary driver were sufficiently far from zero (accordingly, the resonant points are far from the neutral point) and a pair of anti-symmetric resonances, of width $\delta\sim S^{-1/3}$, are clearly defined in that case. As these poles are brought towards $\omega_{0}=0$ (the resonant points approach the neutral point $x=0$) in the small frequency limit, the widths of the resonant layers overlap across $x=0$ by coupling to the reconnection mode. The rate of reconnection and its signatures are strongest when the poles overlap entirely, that is for $\omega_{0}=0$.
@@
Finally, we demonstrated that our solution recovers known scalings of the resonant layer $\delta$ as a function of the Lundquist number $S$ under specific limits. Additionally, we demonstrated new scalings of the reconnection rate and reconnected flux as functions of the driving frequency of the boundary condition. In Figure 5 we show that, in the constant-psi regime, the maximum reconnection rate and reconnected flux are largest for frequencies less than $\omega_{0}\lesssim S^{-3/5}$ and is reduced at larger frequencies until reconnection is completely suppressed for $\omega_{0}\gg S^{-1/3}$.
@@
We conclude that forced reconnection is indeed a particular case of Alfvén resonant absorption and there is a transition from Alfvén resonance to forced reconnection when the forcing frequency $\omega_{0}\ll S^{-1/3}$. However, the reconnection rate and reconnected flux are strongly reduced for $\omega_{0}\gg S^{-3/5}$. When the forcing frequency is $\omega_{0}\gg S^{-1/3}$, Alfvén resonance decouples from the reconnecting mode and the formation of resonant layers away from the neutral line effectively shields reconnection.
Here we have studied the interaction between Alfvén resonances and forced reconnection by considering an idealized symmetric forcing away from the neutral line. However, our approach is general and provides the first step for further investigations of resonances in diverse environments and within different plasma models. In particular, more realistic perturbations can be considered, such as asymmetric and impulsive. This type of forcing is especially of interest for reconnection onset at Earth’s magnetopause. In this regard, recent observations from the MMS spacecraft within Earth’s magnetosphere [27] have identified reconnection events taking place at electron scales, termed electron-only reconnection [28, 29]. By building on prior investigations into forced reconnection within the EMHD framework [30], our approach can be extended to investigate the effect of a time-dependent forcing on electron-only reconnection, with applications to reconnection onset within Earth’s magnetosphere. Additionally, our unified approach holds potential implications for studying the transition from wave-based to reconnection-based coronal heating scenarios [31, 14], which will be explored in future work.
@@ Appendix A Numerical Simulations
$\frac{\partial\psi}{\partial t}=x\frac{d\phi}{dy}+\frac{1}{S}\nabla^{2}\psi$ (55)
$\frac{\partial\Omega}{\partial t}=-x\frac{d}{dy}\nabla^{2}\psi$ (56)
where $\Omega=-\nabla^{2}\phi$ is the vorticity. We impose a boundary condition at $x=\pm 1$ with small amplitude given by:
$\psi(x=\pm 1,y,t)=\Xi_{0}\cos(k_{0}y)\cos(\omega_{0}t)\left[1-\left(1+\frac{t}{\tau}\right)e^{-t/\tau}\right],$ (57)
$\phi(x=\pm 1,y,t)=\pm\Xi_{0}\frac{\sin(k_{0}y)}{k_{0}}\left\{\left(\frac{t}{\tau}\right)^{2}e^{-t/\tau}\cos(\omega_{0}t)-\omega_{0}\sin(\omega_{0}t)\left[1-\left(1+\frac{t}{\tau}\right)e^{-t/\tau}\right]\right\},$ (58)
with $\tau$ being a small growth parameter, following [3]. The code is periodic in the $y$ direction, where the Fast Fourier Transform is used to calculate spatial derivatives. The $x$ direction is non-periodic and a finite difference scheme of the sixth order is used to calculate spatial derivatives [32]. We use an explicit Runge-Kutta of the third order to advance in time. We impose the 2/3 rule for dealiasing with respect to the periodic variable, and a pseudo-spectral filter in $x$ [32]. Simulations are performed by using a grid $L_{x}\times L_{y}=[-1,1]\times[0,2\pi]$ with mesh points $N_{x}\times N_{y}=512\times 32$ for $S=10^{4}$, $N_{x}\times N_{y}=1024\times 32$ for $S=10^{5}$, $N_{x}\times N_{y}=2048\times 16$ for $S=10^{6}$, and $N_{x}\times N_{y}=4096\times 16$ for $S=10^{7}$, with $N_{x}$ chosen to resolve their respective boundary layers. In all simulations we have fixed $\tau=0.001$, $\Xi_{0}=0.001$, and wavenumber $k_{0}=1$ while varying $S$ and the frequency $\omega_{0}$.